# Mitigating Semantic Collapse in Generative Personalization with a Surprisingly Simple Test-Time Embedding Adjustment

### v1

@@ 5.1 Experimental Setup
Metrics. In addition to the CLIP-T and CLIP-I alignment scores introduced in Section 3.2, we also use the DINO image-image alignment score [3] to evaluate the alignment between the generated images and the reference images.
Table 2: Improvement of our TEA over its baselines counterpart.
Method $\text{CLIP}_{T}^{p}$ $\text{CLIP}_{T}^{f}$ CLIP-I DINO-I
CC101
TI+TEA 0.55 0.64 -2.34 -3.81
DB+TEA 0.77 1.87 4.57 0.59
CD+TEA -0.13 0.34 0.37 0.24
CelebA
TI+TEA 0.33 0.57 -2.41 -1.78
DB+TEA 0.51 0.64 -2.37 -2.27
CD+TEA -0.12 0.09 1.56 3.09
5.2 The Effect of Hyper-parameters
@@
Table 1 shows the performance over a range of $\alpha$ values with fixed $\beta=1.5$. Unlike the scaling parameter, the rotation factor $\alpha$ is more sensitive and significantly impacts prompt alignment capability. Increasing $\alpha$ generally improves the model’s ability to capture context, as reflected in the $\text{CLIP}_{T}^{f}$ score, which increases from 26.5 to 28.2 (an improvement of 1.7), and the $\text{CLIP}_{T}^{p}$ score, which rises from 21.4 to 21.9 (a gain of 0.5). While this comes at a minor cost to visual fidelity, with the CLIP-I score dropping from 73.5 to 71.1 when $\alpha=0.2$, the generated images still maintain high visual quality, as shown in Figures 16 and 17.
@@
5.3 Evaluation Results
We conducted an extensive evaluation to validate the effectiveness of the proposed method (Test-time Embedding Adjustment, TEA) when combined with pre-trained personalization methods such as Textual Inversion (TI) [9], DreamBooth (DB) [25], and CustomDiffusion (CD) [16].
The results in Table 6 show that all baselines ‘enjoy’ a significant improvement on prompt alignment when combined with our ‘TEA’, with the average $\text{CLIP}_{T}^{f}$ score increase of 0.54 for TI with TEA, 0.33 for CD with TEA, and 1.9 for DB with TEA, demonstrating substantial gains in text-image alignment. Notably, this improvement is consistent across most concepts as shown in Figure 7, indicating that our method is robust to diverse personalization scenarios, even with fixed, easily chosen hyper-parameters. In terms of visual quality, our method also shows a noticeable boost, with a 4.6 improvement in the CLIP-I score from DB to DB+EA or 0.36 improvement for CD+TEA on the CC101 dataset. Our method also shows interesting effects over the baseline counterparts, as shown in Figure 15, where the SCP in DB lead to the distorted generated images. Our method successfully corrects the semantic embedding and generates more coherent and realistic images. In the CelebA dataset, our method still shows its effectiveness in improving prompt alignment, with a 0.64 improvement in the $\text{CLIP}_{T}^{f}$ score for DB+TEA. There is also a trade-off in image alignment with a 2.37 drop in the CLIP-I score, however, the generated images still maintain high visual quality, as shown in Figures 16 and 17.
Figure 6: Analysis of the effect of scaling factor.
6 Conclusion
To address this, we proposed a training-free test-time embedding adjustment that realigns the personalized embedding with its original semantic context, significantly improving text-image alignment without modifying model weights. Our method, compatible with diverse personalization approaches like Textual Inversion and DreamBooth, demonstrates substantial gains in prompt consistency and image fidelity across varied use cases.
Beyond introducing the SCP problem and providing a practical, straightforward solution, this work also lays the foundation for future research. Potential directions include developing more adaptive approaches that dynamically adjust embeddings based on prompt complexity or context, as well as integrating additional constraints during personalization to better control embedding updates.
Appendix
[…]
Appendix A Related Work
A.1 Diffusion Models
Given a text-to-image diffusion model $\epsilon_{\theta}$, where $\epsilon_{\theta}(x_{t},t,p)$ represents the predicted noise at time step $t$ given the textual embedding $\tau(p)$ of a prompt $p$ and the noisy intermediate vector $x_{t}$ [11, 29, 24], the model is trained by minimizing the following objective:
A.2 Generative Personalization.
Generative personalization task aims to capture personal concepts which are implicitly shared in a reference set of images as generative conditions and then use them as a guided condition to generate new images containing the personal concept. These personal concepts are very difficult to express in the input prompt, e.g., how to express the concept of your dog that is different from a generic dog. Therefore, rather than using prompt engineering techniques to describe the concept in text, this task usually uses a gradient-based method to fine-tune the model parameters to capture the personal concept. There are several categories of generative personalization [2] classified based on the type of generative conditions, such as subject-driven [9, 25, 7, 16, 32], person-driven [33, 31, 8, 6], style-driven [28, 18, 4] or image-driven [23, 34, 35]. Personalizing a T2I diffusion model from only a few examples presents several well-known challenges, including language drift, limited expressiveness of generative conditions, entanglement of concepts, and conditional misalignment.
Textual Inversion and Dreambooth.
While there are many personalization methods have been proposed, they can be traced back to the two representative methods: Textual Inversion (TI) [9] and DreamBooth [25]. Mathematically, given a set of personal images $\mathcal{X}=\{x_{1},x_{2},\cdots,x_{n}\}$ and a pre-trained T2I model $\epsilon_{\theta}$, the goal is to identify a textual embedding $v^{*}$ associated with a specific verbalizable keyword $V^{*}$ (e.g., ‘sks’, ‘<new>’, etc.). This keyword represents the implicit visual concept shared in the reference set $\mathcal{X}$, enabling the model to generate images with the personal concept using any textual prompt $p$ containing the keyword $V^{*}$, e.g., $\lfloor p,V^{*}\rfloor=$ ’A photo of $V^{*}$ playing on a beach’, where .,$\lfloor$.,$\rfloor$ . , . ⌋ is the sentence construction operator.
Textual Inversion (TI) [9] is a pioneering method that proposes obtaining a textual embedding by minimising the following objective:
[…]
A.3 Challenges in Generative Personalization
Language Drift.
Language drift or overfitting occurs due to the limited number of reference images and results in the incorporation of irrelevant elements and the neglect of the textual context within the outputs. Prior works address this issue by introducing preservation mechanisms such as prior-preservation loss [25], locking concept-specific parameters [30] and regularising model weights [10, 21].
Limited Expressiveness of Generative Conditions.
This occurs due to the limited expressiveness of the original textual format and the limited number of tokens allowed in each condition. A common mitigation approach is to use multi-modal conditions such as image-image and sketch-image. To enable pre-trained T2I models to accept new types of conditions and generate in conjunction with the current text prompt conditioning, previous works have attempted to incorporate an additional encoder [36], or add a new adapter module to align internal knowledge of the model with the new condition [20, 13].
Entanglement of Concepts.
The reference image set might include samples that contain both the intended concept and other irrelevant concepts. To effectively isolate and extract the intended concept from the reference set, previous works have employed explicit masks [1, 14, 26] and additional data with the personalized concept [17]. Alternatively, Disenbooth [5] proposed to mitigate the influence of background elements in the reference set by disentangling the identity and background of the reference set.
Conditional Misalignment.
Conditional misalignment is a long-standing problem in the field of conditional generation and is a consequence of the above mentioned challenges, with its cause traced back to the limited alignment power of the original generative model. A largely unaddressed gap in prior personalization work is the potential drift or misalignment of the textual embedding itself during concept learning. We refer this phenomenon as semantic collapse where the learned concept token is still faithful to the visual reference, but fails to retain any meaningful textual semantics and eventually collapses to a simplified form. Existing methods focus on controlling how the model changes through regularisation, but not explicitly ensure that the learned personalized token embedding remains semantically aligned with the context. In this work, we tackle the issue with a lightweight training-free test-time embedding adjustment which is broadly applicable to different personalized methods.
Appendix B Further Quantitative Results
In this section, we provide further quantitative results on the CustomConcept101 dataset and the CelebA dataset. More specifically, we provide the additional results as follows:
  • Additional results showing the semantic collapsing problem in the textual domain measured by different distances in Figure 10.
  • Additional results showing the semantic collapsing problem in images generated by Textual Inversion and DreamBooth in Figures 11 and 12 respectively.
  • Empirical evidence of the semantic shift of the embedding of the entire prompt on DreamBooth in Figure 14.
  • Detailed alignment scores of our TEA over its baselines counterpart on the CS101 and CelebA datasets in Table 6 and Table 7 respectively.
Appendix C Qualitative Results
We provide the additional qualitative results as follows:
  • The cool effects of our method over the baseline counterpart in Figure 15, where the SCP in DB lead to the distorted generated images. Our method successfully corrects the semantic embedding and generates more coherent and realistic images.
  • Qualitative results of the proposed method over its baselines counterpart on the CS101 and CelebA datasets in Figures 17 and 16 respectively.
  • Illustration of the SCP on multiple concepts setting in Figures 19 and 20. The alignment scores are shown in Figure 18.
Appendix D Limitations and Future Work
Limitations on Evaluation.
While we tried our best to design a comprehensive set of experiments to evaluate the proposed method and baselines, there are still some limitations that we would like to address in the future. First, the model used in this paper is the Stable Diffusion v1.5 model, which is one of the most popular pre-trained models in this field. While the proposed method is general and can be applied to other models, we did not test it on larger models, e.g., SDXL or SD version 2.0. Second, we chose three baselines, including Textual Inversion [9], DreamBooth [25], and Custom Diffusion [16], which are among the most popular personalization methods in the field. While the proposed method is general and can be applied to other personalization methods, it might be interesting to apply it to other approaches as well.
@@ Appendix D Limitations and Future Work
In Section 5.2, we provide an analysis of the impact of hyper-parameters on the performance of the proposed method. It has been shown that the performance of the proposed method is sensitive to the rotation factor $\alpha$: the larger the $\alpha$, the more the embedding vector is rotated toward the target concept, improving alignment with the generated images but potentially reducing alignment with the visual concept. In this work, we simply use the same hyper-parameters for all settings. We believe that a search algorithm could be applied to find the optimal hyper-parameters for each prompt, with a stop condition based on the desired alignment with the input prompt.
Appendix E Experimental Setting
E.1 Dataset Construction
@@
Sample sentences from both prompt sets are provided in Table 3, and the full dataset, along with all prompt templates, is available in the anonymous repository at https://anonymous.4open.science/r/Embedding-Adjustment.
Table 3: Sample sentences for $P_{V^{*}}$ and $P^{\text{simple}}_{V^{*}}$. Set $P_{c}$ can be constructed by replacing the word $c$ with $V^{*}$ in the prompt of $P$. All the data and prompts can be found in the anonymous repository.
@@
We use a subset of 9 concepts from the CustomConcept101 dataset as in the original paper [16], each of which has 3-15 images, including ‘Barn’, ‘Tortoise plushy’, ‘Teddy-Bear’, ‘Wooden Pot’, ‘Dog’, ‘Cat’, ‘Flower’, ‘Table’, ‘Chair’ subjects. For the human concept, we use a subset of 10 concepts from the CelebA-HQ dataset [19], which includes 10 identities with 10-15 images per subject. Sample images from the CustomConcept101 and CelebA-HQ datasets are shown in Figure 8 and Figure 9, respectively.
@@
Sample prompts are provided in Table 4, and the full dataset, along with all prompt templates, is available in the anonymous repository at https://anonymous.4open.science/r/Embedding-Adjustment.
Table 4: Sample prompts to generate personalized images. Each prompt consists of a main subject $V^{*}$ and a context $p$. All the data and prompts can be found in the anonymous repository.
@@ E.2 Evaluation Metrics
We use the CLIP-T text-image alignment score [22] to evaluate the alignment between the generated images $\hat{x}$ and the prompts. To have a better understanding of which part of the prompt contributes to construct the generated images, we break down each prompt into multiple segments/concepts and calculate the alignment score for each segment/concept, i.e., ‘$\text{CLIP}_{T}^{f}$’/$\text{CLIP}_{T}^{c}$/$\text{CLIP}_{T}^{p}$ denotes the alignment score of the full prompt, the first segment—the personal concept, and the second segment—the context, respectively. We use the CLIP-I image-image alignment score [22] and DINO image-image alignment score [3] to evaluate the alignment between the generated images and the reference images.
@@
To evaluate this, we use two key metrics based on the CLIP model [22]:
@@ E.3 Computational Settings
Figure 16: Qualitative comparison between baseline methods and their TEA variants on the CelebA dataset. Column 1-2: Reference images. Column 3: TI, Column 4: TI+TEA, Column 5: DB, Column 6: DB+TEA. Input prompts are shown below each image while alignment scores are shown on the top. More results can be found in the anonymous repository.
@@
Figure 17: Qualitative comparisons between baseline methods and their TEA variants on the CustomConcept101 dataset. Column 1-2: Reference images. Column 3: TI, Column 4: TI+TEA, Column 5: DB, Column 6: DB+TEA. Input prompts are shown below each image while alignment scores are shown on the top. More results can be found in the anonymous repository.
Figure 18: Alignment scores showing the SCP in multiple-concept personalization settings. The embedding of $V^{*}_{man}$ held fixed while the embedding of $V^{*}{dog}$ is varied over fine-tuning step. See Figures 19 and 20 for the example images.

### v2

@@ C.1 The Effect of Hyper-parameters
Figure 12: Analysis of the effect of scaling factor.
@@
Table 11 shows the performance over a range of $\alpha$ values with fixed $\beta=1.5$. Unlike the scaling parameter, the rotation factor $\alpha$ is more sensitive and significantly impacts prompt alignment capability. Increasing $\alpha$ generally improves the model’s ability to capture context, as reflected in the $\text{CLIP}_{T}^{f}$ score, which increases from 26.5 to 28.2 (an improvement of 1.7), and the $\text{CLIP}_{T}^{p}$ score, which rises from 21.4 to 21.9 (a gain of 0.5). While this comes at a minor cost to visual fidelity, with the CLIP-I score dropping from 73.5 to 71.1 when $\alpha=0.2$, the generated images still maintain high visual quality, as shown in Figures 23 and 24.
@@
C.2 Surprising Impact of TEA on Anti-DreamBooth
In this section, we present additional results highlighting the surprising impact of TEA on Anti-DreamBooth. Figures 13 and 14 analyze the effect of varying the rotation factor $\rho$ and the scaling factor $\alpha$, respectively.
Interestingly, while scaling has shown its effectiveness in mitigating the SCP in standard personalization settings as discussed in Section C.1, when applied to Anti-DreamBooth, it does little to correct the visual distortion in generated images. By contrast, adjusting the rotation factor $\rho$ shows a pronounced effect such that the generated images become substainally less distorted and better aligned with the protected concept $V^{*}$ as shown in Figure 13. This contrast suggests that the semantic misalignment in anti-personalization settings is more sensitive to directional shift than to embedding magnitude, and that rotation-based corrections are inherently more effective than scaling-based adjustments.
We view this as a valuable and unexpected finding that not only deepens our understanding of why Anti-DreamBooth works (by amplifying the semantic collapse), but also points to promising directions for future work such as geometric interventions or adaptive adjustments during the defense process.
Figure 13: Applying TEA on mitigating the Anti-Personalization Effect of DreamBooth by varying the rotation factor $\rho$
Figure 14: Applying TEA on mitigating the Anti-Personalization Effect of DreamBooth by varying the scaling factor $\alpha$
Appendix D Qualitative Results
We provide additional qualitative results to complement the quantitative analysis:
  • Correcting distorted generations. Figure 22 illustrates how SCP in DreamBooth leads to distorted generations, while TEA corrects the semantic embedding to produce more coherent and realistic images. This connects directly to the discussion on TEA’s unexpected impact on Anti-DreamBooth (Section C.2).
  • Cross-dataset comparisons. Figures 24 and 23 compare TEA-augmented variants of Textual Inversion, DreamBooth, and Custom Diffusion against their baselines on CS101 and CelebA. Further comparisons on Subject Control and Relationship tasks are shown in Figures 28, 29, and 30, highlighting TEA’s robustness across frameworks.
  • SCP in multi-concept prompts. Figures 26 and 27 demonstrate SCP when multiple concepts are combined, with corresponding alignment scores provided in Figure 25.
Appendix E Limitations and Future Work
@@
In Section C.1, we provide an analysis of the impact of hyper-parameters on the performance of the proposed method. It has been shown that the performance of the proposed method is sensitive to the rotation factor $\alpha$: the larger the $\alpha$, the more the embedding vector is rotated toward the target concept, improving alignment with the generated images but potentially reducing alignment with the visual concept. In this work, we simply use the same hyper-parameters for all settings. We believe that a search algorithm could be applied to find the optimal hyper-parameters for each prompt, with a stop condition based on the desired alignment with the input prompt.
Appendix F Experimental Setting
F.1 Dataset Construction
@@
Sample sentences from both prompt sets are provided in Table 3, and the full dataset, along with all prompt templates, is available at https://github.com/tuananhbui89/Embedding-Adjustment.
Table 3: Sample sentences for $P_{V^{*}}$ and $P^{\text{simple}}_{V^{*}}$. Set $P_{c}$ can be constructed by replacing the word $c$ with $V^{*}$ in the prompt of $P$. All the data and prompts can be found in the repository.
@@
We use a subset of 9 concepts from the CustomConcept101 dataset as in the original paper (Kumari et al., 2023), each of which has 3-15 images, including ‘Barn’, ‘Tortoise plushy’, ‘Teddy-Bear’, ‘Wooden Pot’, ‘Dog’, ‘Cat’, ‘Flower’, ‘Table’, ‘Chair’ subjects. For the human concept, we use a subset of 10 concepts from the CelebA-HQ dataset (Liu et al., 2015), which includes 10 identities with 10-15 images per subject. Sample images from the CustomConcept101 and CelebA-HQ datasets are shown in Figure 15 and Figure 16, respectively.
@@
Sample prompts are provided in Table 4, and the full dataset, along with all prompt templates, is available at https://github.com/tuananhbui89/Embedding-Adjustment.
Table 4: Sample prompts to generate personalized images. Each prompt consists of a main subject $V^{*}$ and a context $p$. All the data and prompts can be found in the repository.
@@ F.2 Evaluation Metrics
We use the CLIP-T text-image alignment score (Radford et al., 2021) to evaluate the alignment between the generated images $\hat{x}$ and the prompts. To have a better understanding of which part of the prompt contributes to construct the generated images, we break down each prompt into multiple segments/concepts and calculate the alignment score for each segment/concept, i.e., $\text{CLIP}_{T}^{f}$$\text{CLIP}_{T}^{c}$$\text{CLIP}_{T}^{p}$ denotes the alignment score of the full prompt, the first segment—the personal concept, and the second segment—the context, respectively. We use the CLIP-I image-image alignment score (Radford et al., 2021) and DINO image-image alignment score (Caron et al., 2021) to evaluate the alignment between the generated images and the reference images.
@@
To evaluate this, we use two key metrics based on the CLIP model (Radford et al., 2021):
@@ F.3 Computational Settings
Figure 23: Qualitative comparison between baseline methods and their TEA variants on the CelebA dataset. Column 1-2: Reference images. Column 3: TI, Column 4: TI+TEA, Column 5: DB, Column 6: DB+TEA. Input prompts are shown below each image while alignment scores are shown on the top. More results can be found at repository.
@@
Figure 24: Qualitative comparisons between baseline methods and their TEA variants on the CustomConcept101 dataset. Column 1-2: Reference images. Column 3: TI, Column 4: TI+TEA, Column 5: DB, Column 6: DB+TEA. Input prompts are shown below each image while alignment scores are shown on the top. More results can be found at the repository.
Figure 25: Alignment scores showing the SCP in multiple-concept personalization settings. The embedding of $V^{*}_{man}$ held fixed while the embedding of $V^{*}{dog}$ is varied over fine-tuning step. See Figures 26 and 27 for the example images.
@@
(a) Output from EasyControl
(b) Output from EasyControl with TEA
Figure 28: Comparison of the output from EasyControl pipeline with and without TEA with the same prompt: $V^{*}_{man}$ dog in a construction outfit’ and same random seed. EasyControl with TEA significantly improves the prompt fidelity, mitigating the failure cases of the original EasyControl pipeline, such as the dog stands beside a person or holding a gun. More results showing the same improvement of EasyControl with TEA can be found at https://github.com/tuananhbui89/Embedding-Adjustment.
(a) Output from ReVersion
(b) Output from ReVersion with TEA
Figure 29: Comparison of the output from ReVersion pipeline with and without TEA with the same prompt: ‘cat <R> carrot in the garden’ and same random seed. ReVersion with TEA significantly improves the prompt fidelity, mitigating the failure cases of the original ReVersion pipeline, such as the cat is not carved by the carrot or only shows the carrot. More results showing the same improvement of ReVersion with TEA can be found at https://github.com/tuananhbui89/Embedding-Adjustment.
(a) Output from ClassDiffusion
(b) Output from ClassDiffusion with TEA
Figure 30: Comparison of the output from ClassDiffusion pipeline with and without TEA with the same prompt: ‘barn’ and same random seed. ClassDiffusion with TEA significantly improves the prompt fidelity, mitigating the failure cases of the original ClassDiffusion pipeline, such as the image does not contain a barn but only a sunflower field. More results showing the same improvement of ClassDiffusion with TEA can be found at https://github.com/tuananhbui89/Embedding-Adjustment.